Emily want to purchase a computer that sells for $1250. She can either pay the total amount at the time of purchase, or she can agree to pay the store $115 down and $85 a month for 15 months. How much money can she save by paying the total amount at the time of purchase?
Answer by stanbon(48535)

Emily want to purchase a computer that sells for $1250. She can either pay the total amount at the time of purchase, or she can agree to pay the store $115 down and $85 a month for 15 months. How much money can she save by paying the total amount at the time of purchase?
-----
If she pays all at time of purchase: cost = 1250
If she pays over time : cost = 115 + 85*15 = 1390
-------------------------------------------
She can save 1390-1250 = $140 if she pays all at time of purchase.
==================================================
Cheers,
Stan H.